OpenStudy (lgbasallote):

\[\huge \mathcal L \lbrace t^2 \cos kt \rbrace\] i got as far as \[\mathcal L \lbrace t^2 \cos kt \rbrace = \frac{(s^2+k^2)^2 (-2s) - 4s(k^2 - s^2)(k^2 +s^2)}{(s^2 + k^2)^4}\] what do i do next?

OpenStudy (unklerhaukus):

cancel common factors?

OpenStudy (lgbasallote):

i got \[\Large\frac{-2s(3k^2 -s^2)}{(s^2 + k^2)^3}\]

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

that's right? sweet

OpenStudy (lgbasallote):

so wait...that's right? no joke?

OpenStudy (anonymous):

it is ok .i think in the out side it should be -2s^2 not 2s . otherwise it looks fine.

OpenStudy (lgbasallote):

why :O

OpenStudy (lgbasallote):

are you sure it's squared o.O

OpenStudy (anonymous):

but with what igba wrote in his post after simplifynig i got -2s(3k^2-s^2)

OpenStudy (anonymous):

yes i am sure. what your book says ?

OpenStudy (lgbasallote):

none. this is one of those with no answers

OpenStudy (lgbasallote):

\[\mathcal L \lbrace \cos kt \rbrace = \frac{s}{s^2 + k^2}\] right?

OpenStudy (lgbasallote):

if yes then it seems wolfram agrees with me :DDD http://www.wolframalpha.com/input/?i=second+derivative+of+%28s%2F%28s%5E2+%2B+k%5E2%29%29

OpenStudy (anonymous):

yes it is the cos transformation after first derivative you should have \[\Large \frac{k^2-s^2}{(s^2+k^2)^2}\] right ?

OpenStudy (lgbasallote):

weird...wolfram agrees with my second derivative...but it agrees with your laplace...

OpenStudy (anonymous):

sorry guys i did it :P you were right :P

OpenStudy (lgbasallote):

lol why did you do \[\large \mathcal L \lbrace t^2 s \cos kt \rbrace\] that's cheating :p

OpenStudy (anonymous):

i did it on paper as well :P look above i gave you first derivative :)